OpenStudy (anonymous):

Establish this identity: (2asin(θ)cos(θ))² + (acos²(θ) - asin²(θ))² = a²

OpenStudy (akashdeepdeb):

2asin(θ)cos(θ))² + (acos²(θ) - asin²(θ))² = a² a^2.sin^2 2θ + a^2. cos^2 2θ = a^2 a^2(sin^2 2θ + cos^2 2θ) = a^2 a^2(1) = a^2 LHS = RHS! ESTABLISHED!! :D Understood? :D

OpenStudy (akashdeepdeb):

Use the identities sin 2θ = 2.sinθ.coθ And cos 2θ = cos²θ - sin²θ And sin²θ + cos²θ = 1 Got that? :D

OpenStudy (anonymous):

\[\left\{ a \left( 2\sin \theta \cos \theta \right)\right\}^{2}-\left\{ a \left( \cos ^{2}\theta-\sin ^{2 }\theta \right) \right\}^{^{2}}\] \[=a ^{2}\sin ^{2}2\theta+a ^{2}\cos ^{2}2\theta =a^{2}\left( 1 \right)=a ^{2}\]

OpenStudy (anonymous):

Ohhhhhhhhh. Okay. I do got it! :D I forgot the like terms things. Thanks! ^^

OpenStudy (anonymous):

yw

OpenStudy (akashdeepdeb):

:)